Find the equation of the line using the point-slope formula. Write the final equation using the slope-intercept form.
perpendicular to
8y = x − 4
and passes through the point
(−2, 1)

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Find the equation of the line using the point-slope formula. Write the final equation using the slope-intercept form.
perpendicular to
8y = x − 4
and passes through the point
(−2, 1)
-------------
Do it like this one:
=====================
perpendicular to
6y = x − 3 and passes through the point (−4,1)
----------
Find the slope of 6y = x-3
Solve for y:
y = (1/6)x - 1/2
Slope = 1/6
Slopes perpendicular are the negative inverse
m = -6
=========
Use y-y1 = m*(x-x1) where (x1,y1) is the point
y-1 = -6*(x+4) = -6x - 24
y = -6x - 23
====================================
This is basic. It's like 2+2.
Know it. If you don't, things will get a LOT worse.

Answer by ikleyn(52802)

.

For your benefit, there is a lesson in this site
    - Equation for a straight line perpendicular to a given line and passing through a given point
which explains on how to solve such problems in all details.

From this lesson,  learn the subject once and for all.


Also, as a bonus for you,  consider the lessons in this site
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
that are closely related to your problem.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.

You have a happy opportunity to learn the entire subject from one compact and clearly written source.